# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 10 Apr 2026 v3

@@ Abstract
Classical stochastic-approximation analyses treat the covariance of stochastic gradients as an exogenous modeling input. We show that under exchangeable mini-batch sampling this covariance is identified by the sampling mechanism itself: to leading order it is the projected covariance of per-sample gradients. In well-specified likelihood problems this reduces locally to projected Fisher information; for general M-estimation losses the same object is the projected gradient covariance $G^{\star}(\theta)$, which together with the Hessian induces sandwich/Godambe geometry.
This identification—not the subsequent diffusion or Lyapunov machinery, which is classical once the noise matrix is given—is the paper’s main contribution. It endogenizes the diffusion coefficient (with effective temperature $\tau=\eta/b$), determines the stationary covariance via a Lyapunov equation whose inputs are now structurally fixed, and selects the identified statistical geometry as the natural metric for convergence analysis. We prove matching upper and lower bounds of order $\Theta(1/N)$ for risk in this metric under an oracle budget $N$; the lower bound is established first via a van Trees argument in the parametric Fisher setting and then extended to adaptive oracle transcripts under a predictable-information condition and mild conditional likelihood regularity. Translating these bounds into oracle complexity yields $\varepsilon$-stationarity guarantees in the Fisher dual norm that depend on an intrinsic effective dimension $d_{\mathrm{eff}}$ and a statistical condition number $\kappa_{F}$, rather than ambient dimension or Euclidean conditioning. Numerical experiments confirm the Lyapunov predictions at both continuous-time and discrete-time levels and show that scalar temperature matching cannot reproduce directional noise structure.
Keywords: Stochastic gradient descent; Fisher information; Godambe information; information geometry; diffusion approximation; minimax rates; oracle complexity; batch-size control; simulation optimization
@@ 1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
Classical stochastic approximation (SA) theory provides the analytical machinery—diffusion limits, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—for analyzing such recursions once the gradient-noise covariance is specified. We do not claim novelty for that machinery. Our contribution is logically prior: we identify a structural property of the stochastic-gradient oracle that SA typically treats as exogenous, namely that the mini-batch gradient covariance is pinned down as $b^{-1}G^{\star}(\theta)$ under exchangeable sampling, where $G^{\star}(\theta)$ is the projected covariance of per-sample gradients. In likelihood settings this matrix reduces to Fisher information at the correctly specified point $\theta^{\star}$; for general losses it is the projected gradient covariance that, together with the Hessian $H^{\star}$, induces sandwich/Godambe geometry. Once $G^{\star}(\theta)$ is identified, the diffusion approximation for constant-step SGD is determined rather than modeled; the natural metric for rates is statistical rather than Euclidean; and batch size admits principled design rules under a fixed sampling budget. The subsequent Lyapunov analysis, rate bounds, and complexity guarantees follow by applying standard tools in this identified geometry.
@@ 1.1 Motivation: Simulation Optimization as Sampling Control
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar. Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients. For correctly specified likelihood objectives, $G^{*}(\theta^{\star})$ reduces to projected Fisher information at the well-specified point; for general losses $G^{*}(\theta)$ is the projected gradient covariance that, together with $H^{\star}$, induces sandwich/Godambe geometry. This provides the statistical foundation for interpreting SGD as a controlled diffusion whose effective temperature is $\tau=\eta/b$.
@@
This sampling-control perspective is not limited to simulation optimization. Three canonical MOR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding. The theory developed below applies uniformly to all three classes.
1.2 Core Insights
Before turning to formal statements, we summarize the main idea informally. When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not generic randomness. Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process.
This has three immediate consequences:
  1. 1.
    Mini-batch noise is anisotropic. Doubling the batch size reduces noise equally in all directions, but the shape of the noise ellipsoid is fixed by the identified covariance $G^{\star}$, with Fisher information recovered locally under correct specification.
  2. 2.
    The natural metric for local analysis is statistical, not Euclidean. Since the noise covariance is shaped by $G^{\star}$ near $\theta^{\star}$, measuring convergence in the corresponding statistical metric gives tighter and more interpretable bounds than the Euclidean metric.
  3. 3.
    Batch size is a design variable, not just a hyperparameter. Under a fixed oracle budget $N=Tb$, batch size controls the effective temperature $\tau=\eta/b$ of the induced diffusion. It governs the stationary covariance through the Lyapunov balance, making $b$ a design lever for effort allocation.
The remainder of this section makes these claims precise.
1.3 What We Show
(i) Identification from sampling theory.
Under exchangeable sampling (and simple random mini-batching), the mini-batch gradient satisfies
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mathcal{F}\bigr)\;=\;\frac{1}{b}\,G^{*}(\theta)\;+\;o(b^{-1}),$ (3)
where $G^{*}(\theta)$ is the projected gradient covariance (Section 4). For correctly specified likelihood losses, $G^{*}(\theta^{\star})=F^{\star}(\theta^{\star})$.
(ii) Diffusion and Lyapunov structure.
In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose volatility is determined by the identified covariance $G^{*}(\theta)$ (Section 5). Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}\;=\;\tau\,G^{*}(\theta^{\star}),$ (4)
making explicit that steady-state fluctuations are governed by the ratio of curvature to identified noise.
(iii) Matching rates and minimax lower bounds.
We analyze SGD performance in the frozen statistical geometry at $\theta^{\star}$. We establish $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds via van Trees, showing that the rate and its dependence on intrinsic geometry are tight within the stated model class (Sections 6 and 7).
(iv) Oracle complexity in statistical conditioning.
The oracle complexity bounds scale with a statistical condition number and an effective dimension, isolating statistical difficulty from Euclidean conditioning. This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
(v) Batch size as an OR design variable.
The framework above connects batch-size design to the identified statistical geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
@@ 1.3 What We Show
Theorem 4.3 — Identification of mini-batch covariance under exchangeable sampling: $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)+o(b^{-1})$. Theorem 5.4 — Diffusion approximation with volatility determined by $G^{\star}$. Thm. 5.6, Cor. 5.7 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation. Thm. 6.5, Prop. 6.10 — Matching $1/(Tb)$ upper and lower bounds in the frozen Fisher metric. Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
@@ 1.4 Organization
Section 2 reviews related work. Section 3 defines the oracle model and geometric objects. Section 4 derives the covariance identification from exchangeability. Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge. Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity. Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
@@ 2 Literature Review and Positioning
A substantial literature interprets constant-step SGD as a discretization of an SDE with temperature $\tau=\eta/b$ and various proposed diffusion matrices $\mathbf{C}(\theta)$ (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching. Our covariance identification removes this modeling degree of freedom: under exchangeable mini-batching, the diffusion matrix is determined as $G^{\star}(\theta)$ (Theorem 5.4), and the local OU regime inherits its Lyapunov balance at temperature $\tau$ (Corollary 5.7). Because the identification is logically prior to the discretization corrections of Li et al. (2017) or the Hessian-scaling regime of Smith et al. (2021), it is what enables the subsequent rate-matching bounds.
@@
Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2024). Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
@@
Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives. We state both upper and lower bounds in the identified statistical metric via a van Trees argument (Theorem 6.5 and Proposition 6.10), with complexity governed by statistical curvature and $d_{\operatorname{eff}}$ rather than ambient dimension.
@@
Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient determined by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.7
@@
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.10
@@ 3.2 Data model, objective, and stochastic-gradient oracle
We observe a dataset $z_{1:n}:=(z_{1},\dots,z_{n})$ drawn i.i.d. from $Q$ (or, in the exchangeability formulation of Section 4, conditionally i.i.d. given a directing measure). All lower bounds in Section 6 are stated for a general oracle transcript with bounded predictable information (Assumption 6.9); the i.i.d. sampling model is recovered as a special case. At iteration $t$, the algorithm selects a mini-batch $B_{t}\subseteq\{1,\dots,n\}$ of size $b_{t}$ uniformly without replacement and forms the mini-batch gradient
@@ 3.5 Projection onto identifiable directions
Remark 3.6 (Terminological convention: “Fisher/Godambe”).
Throughout the paper we write “Fisher/Godambe” as shorthand for the following identification: in the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (Fisher information); for general losses, $G^{\star}(\theta^{\star})$ is the projected gradient covariance whose inverse is the Godambe information. Every theorem statement holds with $G^{\star}$ in the general case and specializes to $F^{\star}$ under correct specification. When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding Godambe/sandwich object.

### 15 Apr 2026 v4

# Mini-Batch Covariance, Diffusion Limits, and Oracle Complexity in Stochastic Gradient Descent: A Sampling-Design Perspective

@@ Abstract
Stochastic gradient descent (SGD) is central to simulation optimization, stochastic programming, and online M-estimation, where sampling effort is a decision variable. We study the mini-batch gradient noise as a sampling-design object. Under exchangeable fresh-sampling mini-batches, the conditional covariance given the de Finetti directing measure $\mu$ is $b^{-1}G_{\mu}(\theta)$, and under identifiability the projected population object is $b^{-1}G^{\star}(\theta)$—projected Fisher information for correctly specified likelihoods, the sandwich partner of the Hessian otherwise. This identification fixes the noise matrix entering the diffusion analysis of constant-step SGD: the raw iterate path has a deterministic fluid limit, and the $\sqrt{b/\eta}$-scaled fluctuations satisfy a functional CLT with noise covariance $G^{\star}$; near a nondegenerate optimum the limit is Ornstein–Uhlenbeck, and its Lyapunov covariance scaled by $\eta/b$ matches the linearized discrete recursion at leading order. Under a curvature–noise compatibility condition $\mu_{F}>0$, we prove $1/N$ mean-square upper bounds and an i.i.d. parametric Fisher van Trees lower bound of the same rate order, with oracle-complexity guarantees depending on an effective dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$. Numerical experiments verify the identification and confirm the Lyapunov predictions in direct SGD.
Keywords: Stochastic gradient descent; simulation optimization; sampling-design; mini-batch covariance; diffusion limits; Ornstein–Uhlenbeck approximation; Fisher and Godambe information; oracle complexity; batch-size control
@@ 1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and the linearized local equilibrium of plain SGD exhibits directional structure related to the local curvature–noise geometry, despite the drift being unpreconditioned.
Classical stochastic approximation and diffusion-approximation analyses—small-step SDE surrogates, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—are well-developed OR tools for analyzing such recursions once the gradient-noise covariance is specified (Kushner and Yin, 2003; Harrison, 1985; Whitt, 2002). We do not reprove that machinery. Our contribution is a sampling-design reinterpretation of the noise input: under exchangeable fresh-sampling mini-batches, the conditional mini-batch covariance is determined by the sampling mechanism at $b^{-1}$ times the per-sample gradient covariance, a statement that is elementary but carries structural consequences for operational decisions about sampling budgets. Writing the identified matrix as $G^{\star}(\theta)$—projected Fisher information at the correctly specified point, and more generally the sandwich partner of the Hessian $H^{\star}$—lets us analyze constant-step SGD in a fixed statistical metric rather than a Euclidean one, with the diffusion coefficient determined by the sampling design rather than postulated. Under a curvature–noise compatibility condition $\mu_{F}>0$ (Assumption 6.2; not implied by $H^{\star},F^{\star}\succ 0$ alone), the resulting local rates in the Fisher case match an i.i.d. van Trees lower bound in rate order, and batch size becomes a statistical design variable subject to a sampling budget.
@@ 1.1 Motivation: Simulation Optimization as Sampling Control
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar but a matrix-valued covariance that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. The same sampling-control perspective applies to three canonical MOR problem classes: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is typically small relative to $d$, and the question “how many samples per iterate?” is operationally binding.
1.2 Contributions and scope
Under exchangeable fresh-sampling mini-batches, the conditional mini-batch covariance given the de Finetti directing measure $\mu$ is $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=b^{-1}G_{\mu}(\theta)$ (Theorem 4.3). Under identifiability with a single data-generating distribution, the unconditional covariance reads
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\bigr)\;=\;\frac{1}{b}\,G(\theta),$ (3)
and, after projection onto the identifiable tangent space, we denote the resulting population object $G^{\star}(\theta):=\Pi_{\theta}G(\theta)\Pi_{\theta}^{\top}$ (Section 4). All downstream results are stated in terms of $G^{\star}$. For correctly specified likelihood losses, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$; the finite-dataset-reuse case adds a (1b/n)(1-b/n) correction and is treated separately (Lemma 4.1, Proposition 4.5).
Once $G^{*}(\theta)$ is identified, three consequences follow within the classical stochastic-approximation toolkit. (a) At fixed batch size, constant-step SGD has a deterministic fluid limit (the gradient-flow ODE), and its $\sqrt{b/\eta}$-scaled fluctuations obey a functional CLT converging to a linear SDE with noise covariance $G^{\star}$; near a nondegenerate optimum this is an OU process, and its stationary covariance, multiplied back by $\eta/b$, is the discrete-time Lyapunov benchmark of Proposition 5.13 (Section 5). (b) Mean-square error in the Fisher/identified metric admits a $1/N$ upper bound with an intrinsic effective dimension $d_{\mathrm{eff}}$ and statistical condition number $\kappa_{F}$ (Section 6), matched in rate order by an i.i.d. parametric Fisher van Trees lower bound (Proposition 6.8); geometry dependence appears only in the upper bound. (c) A high-probability oracle-complexity statement follows, with a squared-gradient stationarity rate of order $N^{-1/2}$ (Theorem 7.3).
@@ 1.2 Contributions and scope
Theorem 4.3 — Covariance identification under exchangeable fresh sampling: $\mathrm{Cov}(g_{B}\mid\mu)=b^{-1}G_{\mu}(\theta)$; under identifiability and after projection onto the identifiable tangent space, the population covariance is $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)$. Theorem 5.7 — Fluid limit for the raw iterates and functional CLT for $\sqrt{b/\eta}$-scaled fluctuations, with noise covariance determined by $G^{\star}$. Thm. 5.10, Cor. 5.11 — OU equilibrium at $\theta^{\star}$ on the fluctuation scale; linearized discrete-recursion stationary covariance on the original scale, $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ (Prop. 5.13). Thm. 6.5, Prop. 6.8 — $1/(Tb)$ upper bound and rate-order-matching i.i.d. Fisher lower bound. Theorem 7.3 — Squared-gradient stationarity at rate $N^{-1/2}$; norm stationarity at $N^{-1/4}$.
@@ 1.3 Organization
Section 2 reviews related work. Section 3 defines the oracle model and geometric objects. Section 4 derives the covariance identification from exchangeability. Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge. Section 6 proves an upper bound and a rate-order-matching i.i.d. Fisher lower bound, and Section 7 states oracle complexity. Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
@@ 2 Literature Review and Positioning
A substantial literature interprets constant-step SGD as a discretization of an SDE with temperature $\tau=\eta/b$ and various proposed diffusion matrices $\mathbf{C}(\theta)$ (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching. Our covariance identification removes this modeling degree of freedom. At fixed batch size, we do not claim a nondegenerate diffusion limit of the raw iterate path—its martingale variation is of order $\eta/b$ and vanishes—but we prove a fluid limit plus a functional CLT on the $\sqrt{b/\eta}$-rescaled fluctuations, with noise covariance $C^{\star}{C^{\star}}^{\top}=G^{\star}$ (Theorem 5.7). The local OU regime and the Lyapunov balance then appear as the equilibrium specialization on the fluctuation scale, with $\tau=\eta/b$ recovered on the original iterate scale (Theorem 5.10, Corollary 5.11). This is the standard stochastic-approximation functional CLT architecture (Kushner and Yin, 2003) applied to the fresh-sampling covariance identified in Section 4, and is what fixes the noise input to the subsequent rate analysis.
@@
Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2025). Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
@@
Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives. We give an upper bound in the identified statistical metric that surfaces dependence on $\kappa_{F}$ and $d_{\operatorname{eff}}$ (Theorem 6.5), and match its $1/N$ rate with an i.i.d. van Trees lower bound in the parametric Fisher setting (Proposition 6.8); the lower bound is at the level of rate order and does not certify the geometry dependence.
@@
Continuous-time limit Postulated SDE with ad hoc $\mathbf{C}(\theta)$ Fluid limit (ODE) for raw iterates; functional CLT for $\sqrt{b/\eta}$-scaled fluctuations with diffusion $G^{\star}(\theta)$ Sec. 5; Thm. 5.7
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU fluctuation limit at $\theta^{\star}$; Lyapunov balance $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$ on fluctuation scale; linearized original-scale benchmark $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ (Prop. 5.13 / Cor. 5.11) Sec. 5; Cor. 5.11
@@
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.8
@@ 3.2 Data model, objective, and stochastic-gradient oracle
We observe a dataset $z_{1:n}:=(z_{1},\dots,z_{n})$ drawn i.i.d. from $Q$ (or, in the exchangeability formulation of Section 4, conditionally i.i.d. given a directing measure). The lower bound in Section 6 is stated for the i.i.d. parametric Fisher setting; adaptive-oracle extensions are discussed in Remark 6.11. At iteration $t$, the algorithm selects a mini-batch $B_{t}\subseteq\{1,\dots,n\}$ of size $b_{t}$ uniformly without replacement and forms the mini-batch gradient
@@ 3.5 Projection onto identifiable directions
Remark 3.6 (Terminological convention).
Throughout the paper, the identified noise matrix is $G^{\star}(\theta)$ (the projected gradient covariance). In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (projected Fisher information). For general losses, $G^{\star}(\theta^{\star})$ is not itself the Godambe information; rather, paired with the Hessian $H^{\star}$ it induces the classical sandwich covariance $H^{\star-1}G^{\star}H^{\star-1}$ and, by inversion, the Godambe information $J_{\mathrm{God}}=H^{\star\top}G^{\star-1}H^{\star}$. Every theorem statement holds with $G^{\star}$ in the general case and specializes to $F^{\star}$ under correct specification. When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding object defined by $G^{\star}$; the paired Godambe/sandwich interpretation always requires $H^{\star}$ as well.